# Federated Large Language Models for Swarm Intelligence: A Survey

### v1

@@
Abstract
Federated learning (FL) offers a compelling framework for training large language models (LLMs) while addressing data privacy and decentralization challenges. This paper surveys recent advancements in the federated learning of large language models, with a particular focus on machine unlearning—a crucial aspect for complying with privacy regulations like the Right to be Forgotten. Machine unlearning in the context of federated LLMs involves systematically and securely removing individual data contributions from the learned model without retraining from scratch. We explore various strategies that enable effective unlearning, such as perturbation techniques, model decomposition, and incremental learning, highlighting their implications for maintaining model performance and data privacy. Furthermore, we examine case studies and experimental results from recent literature to assess the effectiveness and efficiency of these approaches in real-world scenarios. Our survey reveals a growing interest in developing more robust and scalable federated unlearning methods, suggesting a vital area for future research in the intersection of AI ethics and distributed machine learning technologies.
{IEEEkeywords}
Federated Learning (FL), Large Language Models (LLMs), Swarm Intelligence, Efficiency, Pre-trained Models, Privacy and Security
@@ 1 INTRODUCTION
In the rapid evolution of artificial intelligence has led to significant breakthroughs in various domains, particularly with the advent of Large Language Models (LLMs) like GPT [1] and BERT [2]. These models have revolutionized the way machines understand and generate human language, enabling applications ranging from automated customer support to advanced natural language processing tasks [3, 4]. Concurrently, there has been a growing interest in swarm intelligence, a field inspired by the collective behavior of decentralized, self-organized systems found in nature, such as ant colonies or bee swarms [5]. Swarm intelligence leverages the simple rules followed by individual agents interacting locally with their environment to solve complex problems, a concept that is inherently robust and scalable [6].
The convergence of LLMs with the principles of swarm intelligence introduces a novel approach to enhancing decentralized decision-making processes [7]. By integrating these two technologies, it becomes possible to process and generate language-based data under the constraints of decentralized and privacy-preserving environments [8]. This integration is especially pertinent in scenarios where centralizing sensitive data is impractical or undesirable, such as in healthcare or financial services, where privacy concerns are paramount.
Federated learning (FL) emerges as a critical enabler in this context. It allows for the collaborative training of LLMs across multiple decentralized nodes without the need to share their actual data [9, 10]. This approach not only preserves the privacy and security of the data but also harnesses the collective intelligence of the swarm. By distributing the learning process across numerous nodes, federated learning enhances the robustness and adaptability of the models, making them more resilient to attacks and capable of generalizing across diverse datasets [11].
However, integrating LLMs with swarm principles in a federated setting introduces unique challenges. These include managing heterogeneous data sources, ensuring consistency in learning outcomes across diverse nodes, and developing efficient communication protocols that adhere to the lightweight nature of swarm agents. Addressing these challenges requires innovative solutions that balance the scalability and flexibility of swarm intelligence with the advanced processing capabilities of LLMs.
Motivated by this, we aim to provide a systematic view of the current research status of federated LLMs for swarm intelligence from various perspectives, including architecture, efficiency, scalability, pre-trained models, security, and privacy. The main contributions of this survey are as follows.
  • Synthesis of Federated Learning and Swarm Intelligence: We provide a systematic review of the methodologies that integrate federated learning with large language models, specifically tailored for swarm intelligence applications. This includes discussions on the adaptations necessary for handling decentralized, autonomous agents in a swarm.
  • Analysis of Techniques and Performance: The paper evaluates various strategies and modifications to traditional federated learning algorithms to enhance their applicability and efficiency in swarm-based systems, emphasizing real-world implementations and case studies.
  • Strategic Insights and Frameworks: We compare a set of criteria and frameworks for designing and deploying federated LLMs within swarm intelligence contexts, offering insights into achieving scalability, robustness, and real-time responsiveness.
The remainder of the survey is as follows. Section 2 provides a detailed exploration of the foundations of swarm intelligence and its integration with large language models, highlighting the synergies and potential enhancements. Section 3 examines the specific adaptations and innovations in federated learning algorithms that facilitate efficient and effective training of LLMs across distributed swarm agents. Section 4 presents a critical analysis of current implementations and discusses the technical challenges, practical limitations, and potential solutions. Section 5 concludes the survey with a discussion on future research directions and the broader implications of federated LLMs for swarm intelligence in various applications.
Figure 1: Federated LLMs for Swarm Intelligence Architecture
2 Foundations and Integration of SI and LLM
2.1 Introduction to Swarm Intelligence
2.1.1 Definition and Historical Context
Swarm intelligence (SI) refers to the collective behavior of decentralized, self-organized systems, typically composed of a population of simple agents interacting locally with their environment and each other [12]. The agents follow very simple rules, and although there is no centralized control structure dictating how individual agents should behave, local interactions between such agents lead to the emergence of “intelligent” global behavior, unknown to the individual agents.
The concept of swarm intelligence originated from biological studies of colonies of ants and swarms of bees in the late 1980s and early 1990s [13]. In nature, these organisms exhibit behaviors that can be extraordinarily complex, such as finding the shortest path to a food source or constructing complex structures, despite the simplicity of individual behaviors and the lack of central control. This paradox, where complex global behavior arises from simple local interactions, has inspired various algorithms and techniques in artificial intelligence.
2.1.2 Key Principles of Swarm Behavior
Swarm behavior is underpinned by several key principles that enable the robust, scalable, and flexible characteristics of natural swarms [14]:
Decentralization: There is no single controlling entity in swarm systems; instead, control is distributed across the individual agents comprising the swarm. This decentralization contributes to the robustness and fault tolerance of the system, as there is no single point of failure.
Simple Rules: Each agent in a swarm follows simple rules, and the interaction of these rules across the entire swarm leads to the emergence of complex behavior. These rules are typically based on the local perception of the environment, without a global view.
Local Interaction: Agents primarily rely on local information or interactions with their neighbors. This limitation is crucial for scalability, as it reduces the communication overhead and computational complexity that would arise from managing global information.
Emergence: The global behavior of the swarm emerges from the interactions of individuals without being explicitly programmed. Emergent behavior is a hallmark of swarm intelligence, providing solutions that are often more adaptive and resilient than those devised through central planning.
Adaptability and Learning: Swarm systems are adaptive; they can change their behavior based on feedback from the environment. This learning can occur at the level of the individual or the swarm, enhancing the system’s ability to respond to dynamic or uncertain environments.
@@ 2.2 Fundamentals of Large Language Models
2.2.1 Architecture and Functionality
Large Language Models (LLMs) are a type of artificial intelligence model designed to understand, generate, and manipulate human language by learning from vast amounts of textual data [15]. These models are based on the architecture of deep neural networks, particularly Transformer [16] models, which have revolutionized the field of natural language processing (NLP) since their introduction.
The core component of a Transformer, and consequently of many LLMs, is the self-attention mechanism. This mechanism allows the models to weigh the importance of different words in a sentence, regardless of their distance from each other in the text [17]. This feature enables the model to capture complex linguistic structures and contextual nuances, making them highly effective for a wide range of NLP tasks, from machine translation and summarization to question answering and text generation.
LLMs such as GPT (Generative Pretrained Transformer) [1], BERT (Bidirectional Encoder Representations from Transformers) [2], and their variants are pretrained on diverse internet-scale datasets using tasks like masked language modeling (where random words are masked out and the model learns to predict them based on the surrounding context) and next-token prediction (predicting the next word in a sequence). This pretraining step allows the models to develop a broad understanding of language and context, which can then be fine-tuned to specific applications with relatively little additional training data [18].
2.2.2 Evolution of LLMs in Natural Language Processing
The evolution of LLMs has marked several milestones in the field of NLP, starting from simpler models like RNNs (Recurrent Neural Networks) [19] and LSTMs (Long Short-Term Memory networks) [20] to more sophisticated architectures like the aforementioned Transformers. The ability of LLMs to handle longer sequences of text and their parallel processing capabilities (a significant advantage over RNNs and LSTMs) have dramatically improved efficiency and scalability in training and inference processes.
The progression from models that could handle only specific tasks (task-specific models) to models that can be applied to virtually any text-based task (general-purpose models) is another crucial development [21]. This shift has been driven by the realization that a deep, contextual understanding of language can be more effectively achieved through large-scale, unsupervised learning from diverse and extensive textual corpora.
As LLMs have grown in size—from millions to billions of parameters—their ability to model and generate human-like text has improved. However, this increase in size comes with its own set of challenges, such as higher computational costs, greater energy consumption, and the need for substantial hardware resources, prompting ongoing research into more efficient model architectures and training methods [22].
2.3 Integrating Swarm Intelligence with LLMs
2.3.1 Potential Benefits and Synergistic Effects
Integrating swarm intelligence (SI) principles with large language models (LLMs) offers unique opportunities to enhance the adaptability and efficiency of distributed computing systems, especially in processing and generating language-based data. This combination leverages the decentralized nature of SI to manage and operate LLMs across multiple nodes, potentially reducing the computational load and data privacy concerns associated with centralized training and inference.
One of the main benefits of this integration is the enhancement of robustness and fault tolerance in language processing tasks. By distributing the LLM’s functionalities across a swarm of agents, the system can continue to operate effectively even if some nodes fail or are compromised, as the collective intelligence of the swarm compensates for individual failures.
Another significant benefit is scalability. Swarm intelligence inherently supports scaling to large numbers of simple agents or nodes, which is particularly useful when deploying LLMs in environments with resource constraints or where real-time responses are crucial. This scalability also allows for more extensive data processing without the need for proportionally increased computational power at a single point.
2.3.2 Case Studies and Successful Implementations
Several case studies highlight the practical applications and benefits of integrating swarm intelligence with large language models:
Decentralized Content Moderation: In social media platforms, a swarm of agents each running smaller segments of an LLM can collaboratively monitor and moderate content in real-time, ensuring compliance with community guidelines while maintaining user privacy.
Multi-Agent Translation Systems: For multinational corporations, swarm-based LLMs can provide real-time translation services across different regions, where each agent handles language tasks specific to its local region, improving the speed and accuracy of services.
Distributed Sentiment Analysis: In market research, a swarm of LLM-equipped agents can perform sentiment analysis on customer feedback data distributed across various servers or locations, synthesizing insights locally and aggregating them to derive global sentiment trends without transferring sensitive data centrally.
3 Federated LLMs for Smarm Intelligence
3.1 LLM Architectures in Federated Settings
Table 1: Overview of Federated Large Language Model Architectures
@@ 3.1 LLM Architectures in Federated Settings
Singhal et al. [24] introduce a novel approach named ”Federated Reconstruction,” which enhances privacy in federated learning by allowing more computation to be performed locally. The paper particularly emphasizes its application to large language models, discussing how it can reduce the amount of data required to be shared with the server, thus enhancing data privacy and reducing bandwidth requirements.
@@
The integration of large language models (LLMs) in federated settings, as explored in these papers, significantly advances the field of swarm intelligence by leveraging the collective learning capabilities of decentralized networks. The first paper’s exploration of N-gram models in federated environments lays the foundational understanding of how individual nodes, like agents in a swarm, can contribute localized knowledge to build a cohesive and comprehensive model. The introduction of [23] in the second paper enhances this concept by minimizing central coordination, thereby promoting more autonomous local decision-making—akin to the independent yet coordinated behavior seen in natural swarms. The  [24] framework further capitalizes on this by scaling up the capabilities of LLMs across diverse and distributed datasets, mirroring how swarms adapt to varied environments. [25] brings these concepts into the realm of edge computing, where LLMs empower edge devices to function like intelligent agents that perform tasks and make decisions in real time, enhancing the swarm’s overall responsiveness and flexibility. Lastly, the techniques discussed in [26] for reducing the parameter count of LLMs ensure that the computational load is manageable even in resource-constrained environments, which is crucial for maintaining the efficiency and agility of a swarm. Together, these studies demonstrate how federated learning models can emulate and enhance swarm-like intelligence, promoting robust, scalable, and decentralized problem-solving capabilities in artificial systems.
@@ 3.2 Efficient Fine-tuning of LLM
Table 2: Summary of Efficient Fine-Tuning of LLMs in Federated Settings
@@ 3.5 Security and Privacy in Federated LLM
Table 5: Security and Privacy in Federated LLM
No. Contributions Technique LM used Dataset Limitations
@@
4 Learned Lessons and Open Challenges
@@
4.1 Learned Lessons
Efficiency in Distributed Environments: Implementing LLMs within federated frameworks has demonstrated significant improvements in computational efficiency. Techniques such as model compression and data pruning have proven effective in reducing the computational load on individual nodes, thereby enhancing overall system efficiency.
Enhanced Scalability: The application of swarm principles has shown that LLMs can be effectively scaled across a larger network of distributed nodes. This scalability is crucial in handling vast amounts of decentralized data, ensuring the LLMs remain effective even as the network expands.
Emergence of Swarm Intelligence: The self-organizing nature of swarm intelligence has been successfully harnessed to improve collective decision-making processes in LLMs. This emergence has been pivotal in enhancing the adaptability of models to new, unforeseen conditions without centralized control.
Improved Privacy Protections: Federated learning inherently supports privacy by design, allowing for the training of LLMs without requiring the sharing of raw data. Advanced cryptographic techniques and secure multi-party computation have further bolstered these protections, ensuring data remains confidential even in collaborative environments.
Robustness and Fault Tolerance: The decentralized approach has inherently improved the robustness and fault tolerance of systems. Swarm-based LLMs are less susceptible to single points of failure, maintaining functionality across the network even when individual nodes encounter issues.

### v2

# 5- BIG DATA MINING AND ANALYTICS

@@
[…]
-5Big Data Mining and Analytics, January 2018, 1(1): $000\headoddnameFirst$ author et al.:Click and Type Your Title Here …
[…]
5- BIG DATA MINING AND ANALYTICS
[…]
5- I S S N 22 2 0 9 6 - 0 6 54   l l 0 ? / ? ? l l p p ? ? ?– ? ? ?
[…]
5- V o l u m e  1 ,   N u m b e r  1,   J a n u a  r y  2 0 1 8
[…]
5- V o l u m e 1, N u m b e r 1, S e p t e m b e lr 2 0 1 8
[…]
3 Recent Advances in Federated Learning Driven Large Language Models: A Survey on Architecture, Performance, and Security
[…]
5 Youyang Qu, Ming Liu, Tianqing Zhu, Longxiang Gao, Shui Yu, and Wanlei Zhou
[…]
Key words:
Federated Learning (FL), Large Language Models (LLMs), Efficiency, Pre-trained Models, Privacy and Security
[…]
6
@@ 1 Introduction
The rapid evolution of artificial intelligence has led to significant breakthroughs across various domains, particularly with the advent of Large Language Models (LLMs) such as GPT [1, 2] and BERT [3, 4]. These models have revolutionized machine understanding and generation of human language, enabling applications ranging from automated customer support to advanced natural language processing tasks [5, 6]. As LLMs continue to scale, the demand for decentralized, privacy-preserving training frameworks has grown, especially in sensitive domains like chemistry [7], healthcare [8], education [9], and finance [10], where data sharing is often constrained by legal and ethical considerations [11].
Federated Learning (FL) has emerged as a promising paradigm that enables collaborative training of LLMs across multiple decentralized clients without transferring raw data [12, 13, 14]. This approach addresses key challenges related to data privacy [15], communication efficiency [16], and compliance with data protection regulations. FL-driven LLMs offer the potential to leverage diverse data sources while preserving individual user privacy, enhancing generalization, and ensuring robustness to data heterogeneity [17, 18, 19].
Despite its promise, the integration of FL with LLMs raises several research challenges. These include handling non-IID (non-independent and identically distributed) data [20], ensuring convergence and consistency across clients [21], minimizing communication overhead, and safeguarding models against privacy attacks such as membership inference and gradient leakage [22]. Moreover, the growing emphasis on data sovereignty has brought forward the need for mechanisms like machine unlearning [23], which allows the removal of individual data contributions from models to comply with policies such as the Right to be Forgotten [24].
Motivated by these emerging needs, this survey aims to provide a comprehensive and structured overview of the current research landscape surrounding Federated Learning for Large Language Models, with an emphasis on architecture, performance optimization, and security and privacy considerations. We highlight recent innovations, practical implementations, and open research challenges at the intersection of federated optimization and large-scale language modeling. The main contributions of this survey are as follows.
  • Comprehensive Overview of Architectures: We present a taxonomy of FL-driven LLM system architectures, outlining design patterns, coordination schemes (centralized, decentralized, hybrid), and training workflows tailored for privacy-preserving large-scale NLP tasks.
  • Analysis of Performance and Adaptability: We assess how FL impacts the training dynamics, convergence speed, and generalization ability of LLMs across heterogeneous environments, and explore optimization techniques such as model compression, personalization, and communication-efficient training.
  • Security and Privacy Perspectives: We review emerging threats to FL-LLMs, such as privacy leakage and poisoning attacks, and survey mitigation techniques including secure aggregation, differential privacy, and machine unlearning strategies.
The remainder of the survey is as follows. Section 2 provides a detailed exploration of the foundations of FL and its integration with LLM, highlighting the synergies and potential enhancements. Section 3 examines the specific adaptations and innovations in FL algorithms that facilitate efficient and effective training of LLMs across distributed agents. Section 4 presents a critical analysis of current implementations and discusses the technical challenges, practical limitations, and potential solutions. Section 5 concludes the survey with a discussion on future research directions and the broader implications of federated LLMs in various emerging domains.
@@
2 Foundations of Federated Learning and Large Language Models
This section provides the foundational background necessary to understand the interplay between Federated Learning (FL) and Large Language Models (LLMs). We first explore the basic principles and motivations of FL, then outline the evolution and architecture of LLMs, and finally describe how the two can be effectively integrated for privacy-preserving and distributed AI solutions.
2.1 Federated Learning: Principles & Motivations
Federated Learning is a decentralized learning paradigm that enables multiple clients, such as edge devices or organizations, to collaboratively train a shared model without exposing their raw data [25, 26]. This is particularly important in domains where privacy, data ownership, or regulatory constraints prevent centralizing datasets [27]. Unlike traditional machine learning pipelines that aggregate data into a central server, FL operates by distributing the learning process [28]. Each client trains a local model on its private data and then shares model updates, such as gradients or weights, with a coordinating server that aggregates them to update the global model [29].
This approach offers several benefits. First, it enhances privacy by keeping data local [30]. Second, it supports scalability across heterogeneous and distributed environments [19]. Third, it provides a viable pathway for compliance with data protection regulations such as GDPR or CCPA [31]. Finally, FL enables local personalization, where clients can adapt the shared model to suit specific needs or user preferences [32]. These features make FL a compelling solution for privacy-sensitive applications in healthcare, finance, mobile computing, and other sectors.
@@ 2.2 Fundamentals of Large Language Models
As FL provides a secure and distributed framework for training, the rise of Large Language Models introduces a new opportunity to apply it to models with unprecedented capability. LLMs, such as GPT [33, 2] and BERT [34, 3], are deep neural architectures trained on vast textual corpora using unsupervised or self-supervised learning objectives. Most modern LLMs rely on the Transformer architecture [35, 36], which introduced the self-attention mechanism that allows models to capture dependencies across arbitrary-length sequences. This design drastically improves upon prior models like RNNs and LSTMs in terms of scalability and contextual understanding [37].
Pretraining tasks such as next-token prediction [38] or masked language modeling [39] enable these models to develop a strong general-purpose understanding of language, which can be fine-tuned with relatively small task-specific datasets. Over the years, LLMs have grown from millions to hundreds of billions of parameters [40], expanding their ability to perform a wide range of language tasks—from summarization and translation to reasoning and code generation [41, 42]. However, this growth comes with challenges: training and deploying LLMs is computationally expensive and typically requires significant infrastructure and engineering effort [43].
2.3 Integration of Federated Learning and LLMs
The convergence of FL and LLMs creates a powerful framework that supports training and deploying large-scale models across data silos without compromising privacy [12]. The motivation for integrating these two technologies stems from the need to utilize sensitive or siloed datasets, such as clinical notes in hospitals [44] or user interactions on mobile devices [45], while adhering to strict privacy constraints. FL provides the decentralized mechanism for training, while LLMs offer a general-purpose, high-capacity architecture capable of adapting to diverse linguistic tasks [46].
Implementing FL for LLMs introduces a number of technical challenges. These include handling heterogeneous data distributions across clients, reducing the communication overhead of model updates, mitigating privacy risks associated with gradient leakage, and achieving consistent convergence despite partial participation [47, 48]. Additionally, there is growing interest in mechanisms like machine unlearning within FL settings, allowing models to forget specific data contributions post-training, which is crucial for compliance with emerging privacy laws [49].
Real-world applications of FL-LLMs are already being explored in sectors such as healthcare, finance, and smart devices. Hospitals can jointly train medical LLMs without sharing patient data, financial institutions can build collaborative assistants while retaining proprietary transaction logs, and edge devices can personalize LLMs locally for user-specific tasks like voice input or text prediction [50, 51]. These scenarios highlight the relevance and promise of FL-LLMs for building secure, personalized, and scalable language-based systems.
[…]
3 Federated Large Language Models
In this section, we present up-to-date research on federated LLMs from the following perspectives: LLM architectures in federated settings, efficient fine-tuning of federated LLM, pre-training of LLM in federated learning, scalable LLM via federated learning, and security & privacy in federated LLM.
3.1 Federated LLM architectures
[…]
@@ 3.1 Federated LLM architectures
Singhal et al. [53] introduce a novel approach named “Federated Reconstruction,” which enhances privacy in federated learning by allowing more computation to be performed locally. The paper particularly emphasizes its application to large language models, discussing how it can reduce the amount of data required to be shared with the server, thus enhancing data privacy and reducing bandwidth requirements.
@@
The integration of large language models (LLMs) in federated settings, as explored in these papers, significantly advances the field of swarm intelligence by leveraging the collective learning capabilities of decentralized networks. The first paper’s exploration of N-gram models in federated environments lays the foundational understanding of how individual nodes, like agents in a swarm, can contribute localized knowledge to build a cohesive and comprehensive model. The introduction of [52] in the second paper enhances this concept by minimizing central coordination, thereby promoting more autonomous local decision-making to the independent yet coordinated behavior seen in natural swarms. The  [53] framework further capitalizes on this by scaling up the capabilities of LLMs across diverse and distributed datasets, mirroring how swarms adapt to varied environments. [54] brings these concepts into the realm of edge computing, where LLMs empower edge devices to function like intelligent agents that perform tasks and make decisions in real time, enhancing the swarm’s overall responsiveness and flexibility. Lastly, the techniques discussed in [55] for reducing the parameter count of LLMs ensure that the computational load is manageable even in resource-constrained environments, which is crucial for maintaining the efficiency and agility of a swarm. Together, these studies demonstrate how federated learning models can emulate and enhance swarm-like intelligence, promoting robust, scalable, and decentralized problem-solving capabilities in artificial systems.
@@ 3.2 Efficient Fine-tuning of LLM
[…]
@@ 3.5 Security and Privacy in Federated LLM
[…]
No. Contributions Technique LM used Dataset Limitations
4 Open Challenges and Future Directions
@@
4.1 Open Challenges and Future Directions in LLM Architectures in Federated Settings
How can we adapt LLM architectures to accommodate diverse computational capacities across federated nodes? Current LLMs are generally designed for centralized, high-performance computing environments, making their deployment in federated settings challenging. One approach is to design adaptive architectures that scale model complexity based on each node’s available resources. For example:
Layer-skipping techniques could allow smaller or less capable devices to bypass certain layers in deep models, focusing on critical layers for inference, while more capable nodes handle the full architecture.
Modular model structures where nodes with limited resources only compute essential components (e.g., embedding layers) while offloading deeper layers to edge or cloud infrastructure. Another approach could involve using lightweight proxy models on resource-constrained nodes to participate in training with minimal computation.
How can we achieve efficient model aggregation without incurring high communication costs in large-scale federated networks? As federated networks scale, the communication costs of aggregating model updates from numerous nodes can be prohibitively high. Developing hierarchical aggregation mechanisms can alleviate this by organizing nodes into tiers, where local aggregations occur before updates are sent to a central server.
Edge-based aggregation tiers could further reduce communication by using intermediary nodes for partial aggregation, reducing the frequency and volume of data sent to the central server.
Quantized and compressed model updates are another promising area. Here, nodes can send reduced precision or partially pruned updates, thus reducing data transfer needs without a significant loss in model accuracy.
How can federated LLMs effectively manage model and data divergence due to non-iid data distributions across nodes? Non-iid data distributions across federated nodes create challenges in training, often leading to biased models. Addressing this requires innovative aggregation and model adjustment techniques:
Adaptive weighting of updates: Develop methods that assign weights to node updates based on similarity to the global data distribution. This could involve clustering nodes based on data profiles, so model updates from similar distributions are aggregated separately before global merging.
Divergence-aware training algorithms: New algorithms could incorporate divergence checks, allowing the model to adapt parameters dynamically based on the node’s data distribution. Exploring approaches like federated multi-task learning where each node’s model partially diverges based on data uniqueness could also help maintain robustness.
How can we design energy-efficient LLM architectures for resource-constrained federated environments? To mitigate the high energy demands of LLMs, future models could focus on parameter-efficient techniques that reduce the computational load:
Sparse and low-rank modeling: Sparse representations or low-rank approximations reduce the number of active model parameters without sacrificing performance. This can be done by pruning layers or enforcing sparsity during training.
Green federated architectures: New architectures could integrate energy monitoring and adaptive energy-use policies that dynamically adjust computation based on each node’s available energy resources, limiting participation in training rounds if necessary to save energy.
Model distillation: Using smaller, distilled versions of large models that retain core functionalities for specific tasks can significantly reduce energy use. Research into task-specific distillation for federated LLMs is an effective way to make LLM deployment feasible on lower-energy devices.
How can federated LLM architectures improve interoperability between edge and cloud infrastructures to leverage their complementary strengths? Designing hybrid architectures that operate seamlessly across edge and cloud infrastructures could enhance the functionality and scalability of federated LLMs.
Hybrid federated models: One approach is to create hybrid models that allow computationally intensive processes to run on the cloud, while edge devices focus on localized or preliminary tasks (e.g., feature extraction). Nodes could switch dynamically between edge and cloud depending on task complexity and resource availability.
Resource-aware task allocation: Develop systems where tasks are automatically assigned based on resource availability, sending large-scale computations to the cloud when edge resources are limited.
Cross-infrastructure synchronization: Federated LLMs could benefit from mechanisms that allow seamless model synchronization between edge and cloud, ensuring consistency and minimizing delays in model updates or inference responses.